2) Use two iterations of the Secant Method to estimate a solution of x + e* - 3 = 0, with 0 and x 1. xo
Answer #1

let fix= x + e ²-3 1.8t Iteration 20=0, x1=1 fixo): fro)= 0+1-3=-2 co f(x) = f(!) = ite-3 = 0.7183 >0 X2 - Xo fixo) (34-80)